What percent of a standard normal distribution is found in each region below? Be sure to draw a graph

a) Z>-1.59

b) -0.4<z<1.8
Answer #1

Part a)


P ( Z > -1.59 ) = 1 - P ( Z < -1.59 )
P ( Z > -1.59 ) = 1 - 0.0559
P ( Z > -1.59 ) = 0.9441

Part b)


P ( -0.4 < Z < 1.8 ) = P ( Z < 1.8 ) - P ( Z < -0.4 )
P ( -0.4 < Z < 1.8 ) = 0.9641 - 0.3446
P ( -0.4 < Z < 1.8 ) = 0.6195